A moving coil instrument have a resistance of 25ohms gives full scale deflection when a current of 20mA is passed through it..Calculate the values of resistance required to enable the instrument to be used as (I) 0-10A ammeter (ii) 0-100v voltmeter


1
Expert's answer
2021-10-04T09:17:21-0400

(i) The resistance is


"r=\frac{r_aI_a}{I-I_a}=\frac{25·20·10^{-3}}{10-20·10^{-3}}=0.05\Omega."

(ii) The resistance is


"r=r_v\bigg(\frac{V}{V_v}-1\bigg)=25\bigg(\frac{100}{25·0.02}-1\bigg)=4975\Omega."